# BOSC: A Backdoor-based Framework for Open Set Synthetic Image Attribution

### v1

@@ Abstract
Synthetic image attribution addresses the problem of tracing back the origin of images produced by generative models. Extensive efforts have been made to explore unique representations of generative models and use them to attribute a synthetic image to the model that produced it. Most of the methods classify the models or the architectures among those in a closed set without considering the possibility that the system is fed with samples produced by unknown architectures. With the continuous progress of AI technology, new generative architectures continuously appear, thus driving the attention of researchers towards the development of tools capable of working in open-set scenarios. In this paper, we propose a framework for open set attribution of synthetic images, named BOSC (Backdoor-based Open Set Classification), that relies on the concept of backdoor attacks to design a classifier with rejection option. BOSC works by purposely injecting class-specific triggers inside a portion of the images in the training set to induce the network to establish a matching between class features and trigger features. The behavior of the trained model with respect to triggered samples is then exploited at test time to perform sample rejection using an ad-hoc score. Experiments show that the proposed method has good performance, always surpassing the state-of-the-art. Robustness against image processing is also very good. Although we designed our method for the task of synthetic image attribution, the proposed framework is a general one and can be used for other image forensic applications.
@@
Synthetic Image Attribution, Backdoor Attacks, Open Set Recognition, Deep Learning for Multimedia Forensics.
@@ I Introduction
The proliferation of generative models, including variational autoencoders (VAEs) [1], generative adversarial networks (GANs) [2, 3], and diffusion models[4], has led to a surge in AI-generated images across the internet, social networks, and various digital environments. In turn, this raised growing concerns about the potential malicious use of such images, for example, in disinformation campaigns or to damage individuals’ reputations.
In response to the threats posed by the diffusion of synthetic and forged images, significant research efforts have been made for the development of synthetic image detectors, capable of distinguishing natural and fake content [5, 6, 7]. In addition to detection, in some cases, it is also important to know which is the specific model or architecture that was used to generate the fake image (hereafter referred to as synthetic image attribution) to gather information about the origin and the synthetic nature of the image. Several methods have been proposed for model-level attribution, that rely on the artefacts or signatures (fingerprints) left by the models in the images they generate [8, 9, 10]. With model-level attribution, models that undergo fine-tuning or retraining with different configurations - such as using different initializations or training data - are regarded as distinct models, since they typically exhibit different fingerprints, reflecting possibly subtle variations introduced during the training process. This can be a limitation in many real-world applications, where model-level granularity is not needed or undesired (e.g. if an attacker steals a copyrighted GAN, and modifies the weights by fine-tuning on a different dataset, model-level attribution is going to fail). Therefore, some recent works have started addressing the attribution task in such a way as to attribute the synthetic images to the source architecture that generated them instead of the specific model [11, 12].
@@
A common drawback of existing attribution methods is their inability to handle images generated by unknown models/architectures, which do not belong to any of the training classes. Such images are wrongly associated with one of the classes used during training, leading to incorrect attributions. In the following, we refer to the classes used during training as in-set classes, while new models/architectures possibly met at test time are termed as out-of-set models/architectures. The difficulty of dealing with out-of-set classes limits the applicability of attribution methods in the real world, where the images analysed at operation time may have been generated by out-of-set models/architectures.
@@
Figure 1: Classification with rejection class: out-of-set images generated by models/architectures that are not included in the list of architectures defined at training time, are classified in an additional rejection class.
@@
A straightforward approach to address the Open-set Synthetic Image Attribution (OSIA) problem is to leverage existing Open-Set Recognition (OSR) methods developed for computer vision applications, that recognize samples from out-of-set classes and do not provide any answer (that would necessarily be wrong) in this case [13, 14, 15]. A schematic representation of such an approach, known as classification with rejection class, is depicted in Fig. 1. Traditional OSR methods rely heavily on the closed-set performance to handle out-of-set classes. Based on the observation that a well-performing classifier usually achieves higher confidence for in-set samples than for out-of-set ones, the confidence of the prediction is used to detect out-of-set samples [16]. The main difference between synthetic image attribution and OSR is that, typically, in computer vision applications, there are clear semantic differences between classes (e.g., cats and dogs), while image attribution relies on subtle statistical artefacts whose use to identify out-of-set samples turns out to be problematic. Thus, deep exploration of the in-set classes may lead to overfitting and reduce closed-set performance for the OSIA problem. Other OSR methods utilize generative models to simulate out-of-set samples, either in the image of the feature domain, creating more compact feature representations for samples belonging to in-set classes, to facilitate the recognition of out-of-set samples [17, 18]. However, simulating realistic out-of-set samples is a tricky problem, for which no really satisfactory solutions exist.
@@
In this paper, we introduce a completely new approach to tackling OSIA. In particular, we introduce a Backdoor-based Open-set Classification (BOSC) framework for multi-class classification with a rejection option, where the concept of backdoor attack [19] is exploited to train a classifier that is able to classify input samples from a certain number of in-set classes, and at the same time, identify samples from unknown classes as out-of-set samples. Backdoor attacks to deep neural networks have first been introduced in [20]. Since then, they have received increasing attention due to the threats they pose to the use of deep neural networks in security-critical applications [19]. A backdoor attack involves injecting a pattern, known as a trigger, into (a subset of) the images used during training, inducing the trained model to exhibit a malicious behaviour, e.g. predicting a wrong predefined class, when fed with images containing the trigger, while continuing to operate as expected on normal inputs. In this work, we exploit the backdoor concept for a benign purpose, namely to create a link between in-set samples and a number of class-specific triggers, and use the absence of such a link to identify out-of-set samples. Specifically, the system we propose works by purposely injecting class-specific triggers inside a portion of the images of the training set to induce the network to establish a connection between class features and trigger features. The behaviour of the trained model in the presence of samples containing the trigger is then exploited at test time to perform sample rejection. In particular, we exploit the fact that no matching trigger can be found for out-of-set samples capable of inducing the backdoor behaviour of the network.
@@
We have verified the effectiveness of BOSC with an extensive set of experiments carried out under various settings. In particular, our experiments show the excellent robustness of BOSC against image post-processing, a property that can be explained by noticing that the trigger is applied to the images under analysis before inputting them to the network, that is, after they have been possibly post-processed. In this way, only the image features are potentially weakened by the processing, while the trigger features remain unaffected. As a result, the matching between the trigger and the class is preserved, and attribution/rejection is performed correctly, even on post-processed synthetic images.
While we designed our method by focusing on open-set synthetic architecture attribution, the approach adopted by BOSC is a general one and can also be used for different image forensic tasks, as demonstrated by the good performance we got by applying it to the classification of synthetic facial attribute editing (see Section VI).
@@
    We propose a new framework for open-set classification that relies on the concept of backdoor attacks and class-specific triggers to train a model for multi-class classification with rejection. Sample rejection is performed based on the output provided by the network when the input samples are tainted with the class-specific triggers.
@@
    We exploit the new framework to develop an open-set classifier for architecture attribution of synthetic images.
@@
    We run several experiments to validate the effectiveness of the classifier in different settings, that is, when different architectures are considered in the in-set pool and for open-set testing. The experiments include comparisons with many state-of-the-art methods, including those specifically proposed for architecture attribution and general algorithms proposed in the ML literature. The experiment also demonstrates: i) the robustness of the proposed method against image post-processing, and ii) the capability of the classifier to deal with unknown models from known architectures [^1] These experiments prove that the system works as it is supposed to do, attributing the images to the source architecture and not to the specific model that was used to generate them..
@@
    We validate the generality of our approach by applying it to a different image forensic task, namely the classification of synthetic facial attribute editing.
@@ II-A Open set recognition
The seminal work on OSR was published by Scheirer et al. in [21], where the authors addressed the problem of determining whether an input belongs to one of the classes used to train a machine learning model or not. In such a work it is shown that, in many cases, simple strategies based on the softmax probabilities or the logit can also effectively judge if a sample comes from an out-of-set class [22], e.g. by exploiting the fact that the maximum output score tends to be smaller for out-of-set inputs [23, 16], or that the energy of the logit vector tends to be lower [24]. Additionally, several works tried to find methods to optimize the representations of in-set and out-of-set samples in the feature space. In [25], Yang et al. designed a suitable embedding space for open set recognition using convolutional prototype learning that removes softmax and implements classification by finding the nearest prototype in the Euclidean norm in the feature space. Multiple prototypes are used to represent different classes. The feature extraction and the prototypes are jointly learned from the data. The method proposed in [26] exploits a different learning framework for OSR, called reciprocal point learning, whereby unknown information is fed to the learner with the concept of reciprocal point to learn more compact and discriminative representations and reduce the risk of misclassifying out-of-set classes as in-set ones.
@@
Another class of works addresses OSR by exploiting the reconstruction error obtained by an Auto-Encoder (AE), assuming that lower reconstruction errors are obtained for in-set classes than for out-of-set ones [27, 28, 29, 14]. For instance, variational autoencoders were exploited in [29], where the authors propose a conditional Gaussian distribution learning framework to detect out-of-set samples by forcing latent features to approximate Gaussian models. Huang et al. [14] proposed to combine the autoencoders with the prototype (PCSSR) and reciprocal learning (RCSSR). Class-specific autoencoders are trained to reconstruct the data based on label conditioning, and the pixel-wise reconstruction errors corresponding to the predicted class, together with semantic-related features, are used for out-of-set sample rejection.
@@
Yet another class of methods is based on adversarial modeling and generative models. Generative adversarial networks (GAN)-based methods resort to GANs to produce unknown-like samples to be used during training. These methods work under the assumption that recognition of out-of-set samples can be achieved by showing to the network a large number of unknown samples during the training process [17, 18, 13, 15]. For instance, G-OpenMax [17] combines the use of generative adversarial networks with the OpenMax method, achieving good performance on the classification of handwritten digits. In ARPL [13] and AKPF [15], reciprocal point learning and prototype learning are enriched by an adversarial mechanism that generates confusing training samples. Specifically, the generated samples are used to optimize the feature space and reduce the so-called open space risk, by restricting the unknown samples in the reciprocal points space ([13]) or learning a kinetic boundary to increase the intra-class compactness and the inter-class separation [15].
@@
Overall, works on OSR mainly improve the open-set performance in three ways: i) developing robust classifiers for in-set samples [25, 16, 30], and using the Maximum Logit Score (MLS) or a similar metric to reject out-of-set samples, ii) thresholding the AE reconstruction error [27, 28, 29, 14], and iii) incorporating synthetic open-set samples in the training process [17, 18, 13, 15]. For the OSIA problem, a good classifier for in-set samples may overfit to known samples and exhibit reduced effectiveness in the open-set case. On the other hand, approaches based on the reconstruction error may be sub-optimal, due to the fact that the fundamental distinction between samples generated by known and unknown generators is based on subtle, visually imperceptible, statistical traces. These traces are often too weak to be effectively thresholded, posing a challenge to methods relying only on the reconstruction error. Finally, reducing the open space risk for synthetic attribution is challenging with a single generator. Generators are designed to produce images with specific semantics and may not be able to generate diverse enough open-set fingerprints.
II-B Synthetic image attribution
@@
Synthetic image attribution has been addressed by relying on both active and passive methods. Active methods inject specific information, e.g., a user-specific key [31] or artificial fingerprints [32, 10], into the generated images during the generation process. These fingerprints or keys are subsequently used during the verification to identify the model. Passive methods rely on the presence within the synthetic images of model-specific, intrinsic artefacts. Passive methods have been proposed in [8, 9, 10, 33, 34, 11, 12]. In particular, Marra et al. [8] first revealed that each GAN leaves a specific fingerprint in the images it generates. The average noise residual image can be taken as a GAN fingerprint. Then, Yu et al. [9] replaced the hand-crafted fingerprint formulation in [8] with a learning-based one, decoupling the GAN fingerprint into a model fingerprint and an image fingerprint.
@@
In addition to model-level attribution, researchers have started proposing approaches that address the attribution problem at the architecture level. In [34], Frank et al. propose to attribute the generated images to the source architecture by relying on the DCT coefficients. The prediction is made in favour of the architecture, which is the most similar to the test image. Yang et al. [11] observed the existence of globally consistent traces across models of the same architecture and proposed an approach to extract architecture-consistent features based on a patchwise contrastive learning framework.
All the above methods focus on closed-set scenarios, limiting the applicability of such methods in real-world applications. To address this limitation, some works have started considering architecture attribution in the open-set scenario. A first step in this direction was made in [35] and [36], where the authors resort to a semi-supervised learning framework that exploits labelled samples from in-set classes and unlabeled samples from out-of-set classes. These samples are used to train a system that, at every step, classifies known samples and clusters the unknown samples, assigning new labels to the new clusters. A drawback of this approach is that the system requires retraining whenever new unknown-class samples are introduced.
@@
Several works have also addressed open-set image attribution by trying to reject out-of-set samples to prevent misclassification [30, 38, 37, 39, 40]. In particular, Wang et al. [30] developed a classifier with a rejection class that exploits a vision transformer-based hybrid CNN architecture with a convolutional and a fully connected branch and performs rejection based on MLS. Fang et al. [38] addressed the OSIA task using a distance-based approach. A test sample is rejected when its minimum distance to the centroids of known in-set classes in the feature space exceeds a predefined threshold. Following a similar approach already proposed for OSR, Yang et al. [37] introduced a progressive open space expansion framework (POSE) to simulate the open space of unknown models through a set of lightweight augmentation models. Abady et al. [40] proposed a verification framework that relies on a Siamese Network, which is capable of determining whether two input images have been produced by the same generative architecture or not. The robust feature representations learnt by the Siamese architecture are also exploited to build a multi-class classifier with rejection.
@@
All these methods strive to enhance the feature representation of the model to learn a compact representation of in-set samples, thereby reducing the open space risk. In this study, we adopt an alternative approach, borrowing the idea of backdoor attacks, to remap the features of multiple classes to a target class position, by injecting class-specific triggers. In this way, we improve the open set performance by relying on the way the system reacts to the presence of the various triggers.

### v2

@@ Abstract
With the continuous progress of AI technology, new generative architectures continuously appear, thus driving the attention of researchers towards the development of synthetic image attribution methods capable of working in open-set scenarios. Existing approaches focus on extracting highly discriminative features for closed-set architectures, increasing the confidence of the prediction when the samples come from closed-set models/architectures, or estimating the distribution of unknown samples, i.e., samples from unknown architectures. In this paper, we propose a novel framework for open set attribution of synthetic images, named BOSC (Backdoor-based Open Set Classification), that relies on backdoor injection to design a classifier with rejection option. BOSC works by deliberately including class-specific triggers inside a portion of the images in the training set to induce the network to establish a matching between in-set class features and trigger features. The behavior of the trained model with respect to samples containing a trigger is then exploited at inference time to perform sample rejection using an ad-hoc score. Experiments show that the proposed method has good performance, always surpassing the state-of-the-art. Robustness against image processing is also very good. Although we designed our method for the task of synthetic image attribution, the proposed framework is a general one and can be used for other image forensic applications.
@@
Synthetic Image Attribution, Backdoor injection, Open Set Recognition, Deep Learning for Multimedia Forensics.
@@ I Introduction
The proliferation of generative models, including variational autoencoders (VAEs) [4], generative adversarial networks (GANs) [5, 6], and Diffusion Models (DM) [7], has led to a surge in AI-generated images across the internet, social networks, and various digital environments. In turn, this raised growing concerns about the potential malicious use of such images, for example, in disinformation campaigns or to damage individuals’ reputations.
@@
In response to the above problems, significant research efforts have been dedicated to the development of tools for tracing back the origin of images produced by generative models, a.k.a. synthetic image attribution. Several methods have been proposed which rely on artifacts or signatures (fingerprints) left by the models in the images they generate [11, 12, 13]. Given that model-level granularity is often not needed or undesired, some recent works have started addressing the attribution task in such a way as to attribute the synthetic images to the source architecture that generated them instead of the specific model [14, 15]. A drawback with most methods developed so far is their inability to handle images generated by unknown models/architectures, which are wrongly associated with one of the classes used during training, leading to incorrect attributions. [^1] In the following, we refer to the classes used during training as in-set classes, while new models/architectures possibly encountered at test time are termed as out-of-set models/architectures. The difficulty of dealing with out-of-set classes limits the applicability of attribution methods in the real world, where the images analyzed at operating time may have been generated by out-of-set models/architectures.
@@
Figure 1: Classification with rejection class: out-of-set class images generated by models/architectures that are not included in the list of architectures defined at training time, are classified in an additional rejection class.
@@
A straightforward approach to addressing the Open-set Synthetic Image Attribution (OSIA) problem involves adapting existing open set classification methods from computer vision, commonly known in machine learning as Open Set Recognition (OSR). These methods identify out-of-set class samples and avoid to classify them (since the classification would necessarily be wrong) [16, 17, 18]. A schematic representation of such an approach, known as classification with rejection class, is depicted in Fig. 1. Traditional OSR methods heavily rely on closed-set performance to handle out-of-set classes. Based on the observation that a well-performing classifier usually achieves higher confidence for samples from in-set classes than out-of-set ones, the confidence of the prediction is used to detect out-of-set classes [19]. For synthetic image attribution however, the deep exploration of in-set class distinguishing features may lead to overfitting and reduce performance for the OSIA problem (while in computer vision applications, there are clear semantic differences between classes - e.g., cats and dogs - image attribution relies on subtle statistical artifacts whose use to identify out-of-set classes turns out to be problematic). Other OSR methods utilize generative models to simulate out-of-set classes, either in the image of the feature domain, creating more compact feature representations for samples belonging to in-set classes, to facilitate the recognition of out-of-set classes [20, 21]. However, simulating realistic out-of-set classes is a tricky problem, for which no really satisfactory solutions exist.
@@
In this paper, we introduce a completely new approach to tackling OSIA, called Backdoor-based Open-set Classification (BOSC), exploiting the backdoor concept to train a classifier that is able to classify input samples from a certain number of in-set classes, and at the same time, identify samples from unknown classes as out-of-set. Backdoor injection into a deep neural network has first been introduced in [23]. Since then, backdoors have received increasing attention due to the threats they pose to the use of deep neural networks in security-critical applications [22]. The integration of a backdoor within a network involves injecting a pattern, known as a trigger, into (a subset of) the images used during training, inducing the trained model to exhibit an anomalous behaviour, e.g. predicting a wrong predefined class, when fed with images containing the trigger, while continuing to operate as expected on normal inputs. In this work, we exploit the backdoor concept for a benign purpose, namely to create a link between samples of the in-set classes and a number of class-specific triggers, and use the absence of such a link to identify samples belonging to out-of-set classes. Specifically, the system we propose works by deliberately injecting class-specific triggers inside a portion of the images of the training set to induce the network to establish a connection between class features and trigger features. The behaviour of the trained model in the presence of samples containing the trigger is then exploited at inference time to perform sample rejection. In particular, we exploit the fact that no trigger capable of inducing the backdoor behaviour can be found for samples belonging to out-of-set classes.
@@
We have verified the effectiveness of BOSC with an extensive set of experiments carried out under various settings, focusing on the prominent domain of face images [2]. In particular, our experiments show the excellent robustness of BOSC against image post-processing, a property that can be explained by noticing that the trigger is applied to the images under analysis before inputting them to the network, that is, after they have been possibly post-processed. In this way, only the image features are potentially weakened by the processing, while the trigger features remain unaffected. As a result, the matching between the trigger and the class is preserved, and attribution/rejection is performed correctly, even on post-processed synthetic images.
While our method was designed for open-set synthetic architecture attribution, the BOSC approach is inherently general and can also be applied to address the open-set scenario in other image forensic tasks. We proved the generality of BOSC by applying it to the classification of synthetic facial attribute editing, as discussed in Section VI. This task involves determining which facial attributes such as age, hair color, expression, or overall physical appearance [1, 2], that have been manipulated by generative models [3].
@@
    We propose a new framework for open-set classification that relies on the concept of backdoor injection and class-specific triggers to train a model for multi-class classification with rejection. Sample rejection is performed based on the output provided by the network when the input samples are tainted with the class-specific triggers.
@@
    We exploit the new framework to develop an open-set classifier for synthetic image attribution. Our experiments validate the effectiveness of the proposed approach by considering a wide variety and types of generative architectures in the in-set and out-of-set, including GANs, DMs and transformers.
@@
    We assessed the robustness of the proposed method against image post-processing - including JPEG compression, Gaussian blur, brightness, contrast and saturation change - and the generalization capability of the classifier, i.e., the capability of the classifier to work well when tested with images generated by unknown models from known architectures [^2] The purpose of these experiments is to prove that the system works as intended, attributing the images to the architecture and not to the specific model that was used to generate them..
@@
    We validate the generality of our approach by applying it to a different open-set image forensic task, namely the classification of synthetic facial attribute editing.
We stress that, both in the case of source attribution and classification of facial attribute editing, we do not aim at distinguishing original and fake/forged content. Rather, our goal is to identify the source of the synthetic images and the kind of manipulation they have undergone.
@@ II-A Open set recognition
The seminal work on OSR was published by Scheirer et al. in [24], where the authors addressed the problem of determining whether an input belongs to one of the classes used to train a machine learning model or not. In such a work it is shown that, in many cases, simple strategies based on the softmax probabilities or the logit can effectively judge if a sample comes from an out-of-set class [25], e.g. by exploiting the fact that the maximum output score tends to be smaller for out-of-set class inputs [26, 19], or that the energy of the logit vector tends to be lower [27]. Starting from Scheirer et al. work, subsequent literature on OSR improved the open-set performance following three approaches: i) developing robust classifiers for samples from in-set classes [30, 31], and using the Maximum Logit Score (MLS) [19] or a similar metric to reject samples from out-of-set classes; ii) performing reconstruction with an Auto-Encoder (AE) and thresholding the reconstruction error [32, 33, 34, 17]; iii) incorporating synthetic open-set samples in the training process [20, 21, 16, 18].
@@
With regard to the first class of methods, techniques to optimize the representations of in-set and out-of-set classes in the feature space were proposed in several works. In particular, in [30], Yang et al. designed a suitable embedding space for open set recognition using convolutional prototype learning that removes softmax and implements classification by finding the nearest prototype in the Euclidean norm in the feature space. Multiple prototypes are used to represent different classes. The feature extraction and the prototypes are jointly learned from the data. A different learning framework is adopted in [31], which proposes to use reciprocal point learning, whereby unknown information is fed to the learner with the concept of reciprocal point to learn more compact and discriminative representations and reduce the risk of misclassifying out-of-set classes as in-set ones.
@@
The second class of works addresses OSR by exploiting the reconstruction error obtained by an AE. These methods work under the assumption that lower reconstruction errors are obtained for in-set classes than for out-of-set ones [32, 33, 34, 17]. Autoencoders were exploited in this sense in [34], where the authors proposed a conditional Gaussian distribution learning framework to detect out-of-set class samples by forcing latent features to approximate Gaussian models. Huang et al. [17] improved the out-of-set performance by combining the use of autoencoders for error reconstruction with the prototype (PCSSR) and reciprocal learning (RCSSR). Class-specific autoencoders are trained to reconstruct the data based on label conditioning, and the pixel-wise reconstruction errors corresponding to the predicted class, together with semantic-related features, are used for out-of-set rejection.
@@
The third class of methods relies on adversarial modeling and generative models to build synthetic open-set samples, which are included in the training process. These methods work under the assumption that recognition of samples from out-of-set classes can be achieved by showing to the network a large number of unknown samples during the training process [20, 21, 16, 18]. Generative adversarial networks (GAN)-based methods resort to GANs to produce unknown-like samples to be used during training. One of these methods is G-OpenMax [20] which combines the use of generative adversarial networks with the OpenMax method, achieving good performance on the classification of handwritten digits. In ARPL [16] and AKPF [18], the adversarial mechanism that generates confusing training samples is used to enrich, respectively, reciprocal point learning and prototype learning, optimizing the features representations for in-set and out-of-set. Specifically, the generated samples are used to optimize the feature space and reduce the so-called open space risk, by restricting the unknown samples in the reciprocal points space ([16]) or learning a kinetic boundary to increase the intra-class compactness and the inter-class separation [18]. Theoretical works have also been proposed addressing the OSIA problem [28, 29]. In particular, in [28], the authors demonstrated that OSR performance is correlated with feature diversity, and that learning diverse representations is crucial in reducing the open space risk.
@@
For the OSIA problem, the off-the-shelf application of the above general OSR methods may not be very effective. A well performing classifier for synthetic image attribution may overfit to known samples and exhibit reduced effectiveness in the open-set case. On the other hand, approaches based on the reconstruction error may be sub-optimal, due to the fact that the fundamental distinction between samples generated by known and unknown generators is based on subtle, visually imperceptible, statistical traces. These traces are often too weak to be effectively thresholded, posing a challenge to methods relying only on the reconstruction error. For the same reason, it is not obvious that relying on a fusion of diverse feature representations that are discriminative for closed-set classes can reduce the open-space risk. Finally, reducing the open space risk for synthetic attribution is challenging with a single generator. Generators are designed to produce images with specific semantics and may not be able to generate diverse enough open-set fingerprints. These limitations have been confirmed in some recent works [35, 43] and have pushed researches to develop specific methods for OSIA.
II-B Synthetic image attribution in closed and open set
@@
Synthetic image attribution has been addressed by relying on both active and passive methods. Active methods inject specific information, e.g., a user-specific key [36] or artificial fingerprints [37, 13], into the generated images during the generation process. These fingerprints or keys are subsequently used during the verification to identify the model. Passive methods rely on the presence within the synthetic images of model-specific, intrinsic artifacts. Passive methods have been proposed in [11, 12, 13, 38, 39, 14, 15]. In particular, Marra et al. [11] first revealed that each GAN leaves a specific fingerprint in the images it generates. The average noise residual image can be taken as a GAN fingerprint. Then, Yu et al. [12] replaced the hand-crafted fingerprint formulation in [11] with a learning-based one, decoupling the GAN fingerprint into a model fingerprint and an image fingerprint.
@@
In addition to model-level attribution, researchers have started proposing approaches that address the attribution problem at the architecture level. In [39], Frank et al. propose to attribute the generated images to the source architecture by relying on the Discrete Cosine Transform (DCT) coefficients. The prediction is made in favor of the architecture which is the most similar to the test image. Yang et al. [14] observed the existence of globally consistent traces across models of the same architecture and proposed an approach to extract architecture-consistent features based on a patchwise contrastive learning framework.
@@
All the above methods focus on closed-set scenarios, limiting the applicability of such methods in real-world applications. To address this limitation, some works have started considering architecture attribution in the open-set scenario. A first step in this direction was made in [40] and [41], where the authors resort to a semi-supervised learning framework that exploits labeled samples from in-set classes and unlabeled samples from out-of-set classes. These samples are used to train a system that, at every step, classifies known samples and clusters the unknown samples, assigning new labels to the new clusters. A drawback of this approach is that the system requires retraining whenever new unknown-class samples are introduced. Several works have also addressed open-set image attribution by trying to reject out-of-set classes to prevent misclassification [35, 43, 42, 44, 45]. In particular, Wang et al. [35] developed a classifier with a rejection class that exploits a vision transformer-based hybrid CNN architecture with a convolutional and a fully connected branch and performs rejection based on MLS. A similar approach to address the OSIA problem, but resorting to a distance-based method for sample rejection, was proposed by Fang et al. [43]. A test sample is rejected when its minimum distance to the centroids of known in-set classes in the feature space exceeds a predefined threshold. Following a similar approach already proposed for OSR, Yang et al. [42] introduced a progressive open space expansion framework (POSE) to simulate the open space of unknown models through a set of lightweight augmentation models. A completely different method, which addresses the OSIA task as a verification problem, is proposed by Abady et al. [45]. In this paper, the authors introduce a verification framework, that relies on a Siamese Network, for determining whether two input images have been produced by the same generative architecture or not. The robust feature representations learned by the Siamese architecture are also exploited to build a multi-class classifier with rejection, using an approach similar to the one in [35].
@@
All these methods strive to enhance the feature representation of the model to learn a compact representation of in-set class samples, thereby reducing the open space risk. In this study, we propose a completely different approach to address the OSIA problem, borrowing the idea of backdoor attacks, to remap the features of multiple classes to a target class position, by injecting class-specific triggers. In this way, compactness is achieved for closed-set samples in the backdoor space, where a small open-space risk can be achieved. Open set samples are then revealed by relying on the way the system reacts to the presence of the various triggers.